# Navigate the Unknown: Enhancing LLM Reasoning with Intrinsic Motivation Guided Exploration

### v3

@@ 2.2.3 Advantage-preserving Reward Implementation
By applying trajectory-aware exploration rewards with dynamic reward scaling and advantage-preserving implementation, i-MENTOR efficiently enhances LLM reasoning by guiding the model toward diverse possible responses during training, while significantly improving overall reasoning ability. The complete algorithm is detailed in Appendix B.
@@ 3 Experiments
  • Q1: How does i-MENTOR enhance LLM reasoning performance in comparison with vanilla PPO and GRPO baselines?
  • Q2: What is the individual contribution of i-MENTOR’s core components to its overall effect?
In the following subsections, we first present the experimental setup, followed by a systematic analysis of results to address these questions. Notably, we also provide training statistics in Figure 3, a case study in Figure 2 and Appendix A, and comparisons with basic exploration techniques in Appendix D.
@@ 3.1.1 Dataset
We conduct experiments on three public datasets, i.e., GSM8K  [^2] https://huggingface.co/datasets/openai/gsm8k[9], Countdown-34  [^3] https://huggingface.co/datasets/Jiayi-Pan/Countdown-Tasks-3to4[11] and its harder version Countdown-4  [^4] https://huggingface.co/datasets/Jiayi-Pan/Countdown-Tasks-4[11] to validate i-MENTOR’s effectiveness against vanilla PPO and GRPO algorithms. For computational efficiency, we use a subset of the complete dataset of Countdown-34 and Countdown-4 for training. The detailed dataset statistic is shown in Appendix C.
@@ 3.1.2 Evaluation Protocol
Environment and LLM backbone: The experiment environment is built on the TinyZero  [^5] https://github.com/Jiayi-Pan/TinyZero[26] and verl  [^6] https://github.com/volcengine/verl framework with Qwen2.5-3B [44, 35] as base model. Similar to DAPO [46], we exclude the KL penalty from all RL algorithms after a detailed analysis in Appendix D.
Network structure of i-MENTOR: The predictor and target networks in i-MENTOR share the same architecture, which is dataset-agnostic and LLM-independent. For computational efficiency, we simply apply a lightweight predictor and target network structure for i-MENTOR (embedding size$=16$, three FFN layers with neurons $[16,8,1]$), ensuring negligible additional training time. To make a fairer comparison, we apply a set of fixed training hyperparameters to all three datasets.
Evaluation metric: For evaluation, we report the average accuracy on five experiments instead of evaluation scores to disable the potential gains brought by different format rewards during evaluation.
Additionally, we include implementation details in Appendix E to further elaborate on the training settings.
@@ 3.2 Main Result (Q1)
The main results are presented in Table 1. Additionally, Figure 4 highlights key training details via some of the training trajectories, showcasing the performance improvements achieved during training.
Table 1: Average accuracy on the three datasets. i-MENTOR-PPO denotes the implementation of i-MENTOR on PPO, while i-MENTOR-GRPO denotes the implementation of i-MENTOR on GRPO. “Improve” indicates relative improvement.
@@
From Table 1 and Figure 4 we could conclude that:
@@
  • A key challenge in improving LLM reasoning lies in handling difficult samples, which typically yield near-zero rewards that hinder parameter updates. Our experiments reveal an interesting pattern: i-MENTOR’s improvements are most significant on Countdown-4, followed by Countdown-34, while on GSM8K, where all RL methods converge within just 20 steps, the improvements are comparatively less pronounced. This progression aligns with the relative difficulty levels of these datasets, where Countdown-4 > Countdown-34 > GSM8K. The results suggest that i-MENTOR’s dense exploration rewards for each response sequence specifically help models overcome learning barriers in challenging samples.
@@
3.3 Ablation Study (Q2)
Table 2: Ablation study on Countdown-34.
Model Accuracy
GRPO 0.6711
$+$Trajectory-aware Exploration Rewards 0.6939
$+$Dynamic Reward Scaling: Exploration Reward Regularization 0.6988
$+$Dynamic Reward Scaling: Error-conditioned Reward Trigger 0.7007
$+$Dynamic Reward Scaling: Policy-preserving Exploration Attenuation 0.7065
$+$Advantage-Preserving Reward Implementation (i-MENTOR-GRPO) 0.7132
From the table, we could conclude that: (1) Incorporating trajectory-aware exploration rewards leads to a notable improvement in reasoning performance, increasing the accuracy from 0.6711 to 0.6939. This demonstrates that enabling the model to explore diverse potential responses during training promotes a richer learning process. By encouraging the discovery of alternative reasoning paths rather than converging prematurely on a limited set of high-reward responses, the model acquires more nuanced reasoning capabilities, as evidenced by the marked accuracy gain. (2) Adding the dynamic reward scaling mechanism further enhances model performance, with accuracy improving from 0.6939 to 0.7065. By applying exploration reward regularization, error-conditioned reward trigger, and policy-preserving exploration attenuation, i-MENTOR conducts efficient exploration with adaptive balance on exploration and exploitation during policy updates, ensuring stable and effective policy optimization. (3) Finally, the advantage-preserving reward implementation ensures seamless compatibility with RL algorithms, such as PPO and GRPO. By preserving the statistical integrity of outcome-based advantage distributions while incorporating exploratory signals, i-MENTOR achieves a robust balance between exploiting known strategies and exploring new areas, thus enhancing generalization capabilities. With all three components, i-MENTOR-GRPO demonstrates the highest improvement in accuracy, raising it to 0.7132.
@@ 4.1 Reinforcement Learning for LLMs
Reinforcement learning has evolved from foundational approaches like PPO [31] to advanced frameworks such as GRPO [33]. PPO achieves stability in policy updates through clipped objectives and trust region constraints, effectively aligning models for dialogue systems and code generation [38, 34]. Its architecture treats token positions in reasoning trajectories of LLM as distinct states for advantage estimation [16], though joint policy-value optimization introduces computational overhead [33].
GRPO [33] mitigates these limitations through trajectory-level sampling and group-wise advantage normalization. By generating multiple responses per input and standardizing rewards across trajectory groups, GRPO reduces policy update bias while maintaining efficiency. This sequence-level evaluation mechanism decouples advantage estimation from computational complexity, proving particularly effective in multi-step reasoning scenarios where traditional RL struggles with action space dimensionality. Recent implementations such as DAPO [46] and Dr. GRPO [21] further showcase GRPO’s scalability through detailed refinement of reasoning objectives, though exploration efficiency remains constrained by static and sparse reward structures [10, 12]. Differently, i-MENTOR enhances LLM reasoning through providing dense, stable, and efficient exploration rewards in RL optimization process. These rewards enable intrinsic motivation-guided exploration over diverse response trajectories during training, systematically improving reasoning capabilities.
@@ 4.2 Improving Reasoning Ability of LLMs
Recent advances in LLM reasoning focus on three key paradigms:
Pre-training augmentation improves foundational reasoning by exposing models to curated mathematical datasets and synthetic traces [45, 13, 33]. Models like Qwen [45] and Llama [13] show significant gains through domain-specific data scaling, though this requires heavy computational resources and careful data curation.
Prompt engineering [29, 7] enhances reasoning via structured inputs. Techniques like Chain-of-Thought [47, 8] decompose problems into steps, while self-consistency [39] improves outputs through majority voting. Extensions, such as multi-agent debates [18] and iterative refinement [23], further extend capabilities but are limited by manual design and high computational costs.
Algorithmic enhancement combines inference-time search with RL-based fine-tuning. Methods like Monte Carlo Tree Search [4, 42], beam search [43], and RL approaches (e.g., PPO [31] and GRPO [33]) optimize reasoning policies but often overexploit fixed high-reward paths, limiting exploration of novel solutions. Unlike these methods, our approach actively incentivizes exploration of under-optimized paths through exploration rewards, enabling deeper action-space traversal.
@@ 5 Conclusion
We present i-MENTOR, a method that addresses sparse rewards and enhances explorations in RL-based LLM reasoning optimization through three key advances. First, our trajectory-aware exploration reward mechanism uses lightweight neural networks to provide dense intrinsic rewards without adding computational overhead or episodic length bias, enabling efficient discovery of high-quality responses. Second, dynamic reward scaling automatically balances exploration and exploitation in large action spaces, ensuring sustained exploration while maintaining policy stability. Third, our advantage-preserving reward implementation decouples exploration incentives from outcome-based gradients, resolving conflicts between intrinsic and extrinsic rewards. Experiments show that i-MENTOR significantly improves reasoning performance with both vanilla PPO and GRPO, demonstrating its versatility and effectiveness. By stabilizing exploration-exploitation dynamics, i-MENTOR enables efficient training and diverse trajectory discovery in complex reasoning tasks.
@@ Appendix A Detailed Case Study
Through the methodological components described in Section 2.2, i-MENTOR implements a reinforcement learning framework with structured exploration guidance to enhance LLM reasoning capabilities. The complete optimization procedure, formalized in Algorithm 1, operates as follows: Given an input batch $B$ of question-response pairs $(q,o)$ for advantage estimation, we first compute the baseline outcome-based reward signals $R$ and initial advantage estimates $\hat{A}_{\text{old}}$ using conventional outcome-based reward functions and RL advantage estimation methods. Subsequently, each sample is processed through i-MENTOR’s dual networks $(f_{\theta_{P}},f_{\theta_{T}})$ to simultaneously update the policy network $f_{\theta_{P}}$ and obtain trajectory-aware exploration rewards $R^{\star}_{2}(q,o)$. These exploration rewards are then adaptively scaled through our dynamic reward scaling approaches, yielding $R^{\star}(q,o)$ that maintain training stability across different optimization phases. Finally, the refined advantages $\hat{A}_{\text{new}}$ are computed through our advantage-preserving implementation mechanism that injects $R^{\star}(q,o)$ into $\hat{A}_{\text{old}}$ without distorting the original policy gradient signals. The resulting $\hat{A}_{\text{new}}$ subsequently drives policy updates through standard RL optimization algorithms like PPO and GRPO, enabling effective LLM optimization while maintaining gradient stability.
@@
4:  Obtain $R^{\star}_{2}(q,o)$ by predicting $r(q,o)$ and conduct min-max scaling in $B$ via Equation (6)
5:  Obtain $R^{\star}_{3}(q,o)$ from error-conditioned reward trigger via Equation (7)
6:  Obtain $R^{\star}(q,o)$ from policy-preserving exploration attenuation via Equation (8)
7:  Add the exploration reward $R^{\star}(q,o)$ to $\hat{A}_{old}$ for a new advantage $\hat{A}_{new}$ via Equation (9)
8:  return $\hat{A}_{new}$
@@ Appendix C Dataset Statistics
Appendix D Comparison with Basic Exploration Techniques
@@
Table 4: Comparison with Other Naive Exploration Methods. “$\beta$” and “Temperature” indicate the KL penalty and LLM temperature coefficients.
@@
[…]
Appendix E Implementation Details
Appendix F Limitations
@@ Appendix G Parameter Sensitivity Analysis
i-MENTOR involves two hyperparameters: $\alpha$ (Equation (6)), which scales the maximum exploration reward intensity, and $\gamma$ (Equation (8)), which regulates its decay rate. In this section, we visualize the parameter sensitivity experiments on Countdown-4 for i-MENTOR-GRPO in Figure 6 to analyze the influence of both on the performance of i-MENTOR.

### v4

@@ 2.2.3 Advantage-preserving Integration
By applying trajectory-aware exploration rewards with error-conditioned reward allocation and advantage-preserving integration, i-MENTOR efficiently enhances LLM reasoning by guiding the model toward diverse possible responses during training, while significantly improving overall reasoning ability. The complete algorithm is detailed in Appendix B.
@@ 3 Experiments
  • Q1: How does i-MENTOR enhance LLM reasoning performance?
  • Q2: How does i-MENTOR generalize across models and data distributions?
In the following subsections, we begin by outlining the experimental setup, followed by a systematic results analysis addressing our core questions. Training statistics are visualized in Figure 3. Appendix A provides an expanded case study for Figure 2, while implementation specifics–including pseudo code (Appendix B), data statistics (Appendix C), and configuration details (Appendix D)–are deferred for completeness. Beyond core experiments, Appendix E validates i-MENTOR’s superiority over basic exploration techniques. We further include an ablation study (Appendix F) quantifying submodule contributions and a sensitivity analysis (Appendix G) examining hyperparameter impacts on performance.
@@ 3.1.1 Dataset
We conduct experiments on three public datasets, i.e., GSM8K(Cobbe et al., 2021), Countdown-34 and its harder version Countdown-4(Gandhi et al., 2024) to validate i-MENTOR’s effectiveness against vanilla PPO and GRPO algorithms. For computational efficiency, we use a subset of the complete dataset of Countdown-34 and Countdown-4 for training. The detailed dataset statistic is shown in Appendix C. Additionally, we also provide detailed experiments on the AIME 2024 dataset between GRPO and i-MENTOR-GRPO to further validate i-MENTOR’s effectiveness with multiple outputs in Table 2.
@@ 3.1.2 Evaluation Protocol
Environment and LLM backbone: The experiment environment is built on the TinyZero  [^2] https://github.com/Jiayi-Pan/TinyZero(Pan et al., 2025) and verl  [^3] https://github.com/volcengine/verl framework with Qwen2.5-3B (Yang et al., 2024a; Team, 2024) as base model. We also provide generalization experiments on deepseek-7b in Section 3.4.1. Similar to DAPO (Yu et al., 2025), we exclude the KL penalty from RL algorithms after a detailed analysis in Appendix E.
Network structure of i-MENTOR: The predictor and target networks in i-MENTOR share the same architecture, which is dataset-agnostic and LLM-independent. For computational efficiency, we simply apply a lightweight predictor and target network structure for i-MENTOR (embedding size=16=16, three FFN layers with neurons $[16,8,1]$), ensuring negligible additional training time. To make a fairer comparison, we apply a set of fixed training hyperparameters to all datasets.
Evaluation metric: For evaluation, we report the average accuracy on five runs for each experiment.
Implementation details can be found in Appendix D to further elaborate on the training settings.
@@ 3.2 Main Result (Q1)
The main results are presented in Table 1. Figure 4 highlights key training details via some of the training trajectories, showcasing the performance improvements achieved during training.
Table 1: Average accuracy. i-MENTOR-PPO/GRPO denotes the implementation of i-MENTOR on PPO/GRPO. “Improve” indicates relative improvement.
@@
From Table 1,  2 and Figure 4 we could conclude that:
@@
  • A key challenge in improving LLM reasoning lies in handling difficult samples, which typically yield near-zero rewards that hinder parameter updates. Our experiments reveal an interesting pattern: i-MENTOR’s improvements are most significant on Countdown-4, followed by Countdown-34 and GSM8K. This progression aligns with the relative difficulty levels of these datasets, where Countdown-4 > Countdown-34 > GSM8K. The results suggest that i-MENTOR’s dense exploration rewards specifically help models overcome learning barriers in challenging samples. This phenomenon represents a further contribution of our study: it originates from RLVR’s core distinction from traditional RL–leveraging a pre-trained LLM instead of a randomly initialized model. Consequently, exploration yields significantly greater benefits when addressing complex problems.
@@
3.3 Analysis on Multiple Generations (Q1)
In this section, we validate i-MENTOR’s efficacy on the AIME 2024 benchmark between GRPO and i-MENTOR-GRPO. For each question, the LLM generates multiple answers to enable more rigorous performance evaluation (Yu et al., 2025). Results in Table 2 demonstrate two key findings: First, LLM accuracy improves consistently with increasing response counts, confirming that aggregating multiple outputs enhances model performance. Second, i-MENTOR achieves approximately 20% improvement across all metrics compared to GRPO baselines. This validates i-MENTOR’s effectiveness in solving complex problems while efficiently leveraging multiple responses.
Table 2: Experimental result on AIME 2024.
Method Pass@1 Pass@8 Pass@32 Avg@8 Avg@32
GRPO 0.3 0.3 0.3333 0.3 0.3073
i-MENTOR-GRPO 0.3667 0.3667 0.4 0.3667 0.3667
Improve 22.23% 22.23% 20.01% 22.23% 19.33%
3.4 Generalization Analysis (Q2)
In this section, we conduct experimental analysis to verify the generalization ability of i-MENTOR from both the model and dataset perspectives.
3.4.1 Base Model Generalization Analysis
Here, we verify the effectiveness of i-MENTOR on a new commonly used base model, deepseek-7b, with GRPO on the GSM8K dataset to validate i-MENTOR’s generalization on different base models. From the result in Table 3, we could conclude that i-MENTOR can be applied to different base models, including Qwen and DeepSeek, and provide performance gains for their reasoning process.
Table 3: Experimental result with deepseek-7b base model on GSM8K.
Dataset GRPO i-MENTOR-GRPO
GSM8K 0.6535 0.6641
3.4.2 Dataset Generalization Analysis
This subsection examines i-MENTOR’s impact on generalization capability by evaluating performance on out-of-distribution data. Specifically, we train the LLM exclusively on Countdown-4 and measure performance on the unseen Countdown-34 dataset. As evidenced by Table 4, i-MENTOR-GRPO surpasses GRPO on both Countdown-4 and the held-out Countdown-34 dataset after training. This demonstrates that i-MENTOR not only enhances optimization efficacy but also preserves - and even strengthens - the model’s generalization capacity when exposed to novel data distributions.
Table 4: Dataset Generalization analysis.
Dataset GRPO iMENTOR-GRPO Improve
Coundown-4 0.3831 0.4726 23.36%
Countdown-34 0.3369 0.3877 15.08%
Table 5: Training time analysis.
Dataset PPO i-MENTOR-PPO GRPO i-MENTOR-GRPO
GSM8K 0.70h 0.73h 0.86h 0.88h
Countdown-34 7.43h 7.81h 12.13h 14.55h
Countdown-4 6.97h 7.20h 11.69h 13.03h
4 Training Time Analysis
Here we summarize the training time for experiments in Table 1 to examine the additional cost in terms of time that i-MENTOR incurs. The results in Table 5 demonstrate that i-MENTOR introduces only minimal training time overhead versus baselines without it. This efficiency stems from the predictor and target networks requiring small parameter counts–they comprise several compact linear layers. These networks compute rewards via output differences to determine trajectory visit frequency during training, without needing complex semantic understanding. Critically, since the exploration module operates solely during training, i-MENTOR has zero inference overhead. This confirms i-MENTOR’s superiority in delivering performance gains with negligible training time impact.
@@ 5.1 Reinforcement Learning for LLMs
Reinforcement learning has evolved from foundational approaches like PPO (Schulman et al., 2017) to advanced frameworks such as GRPO (Shao et al., 2024). PPO achieves stability in policy updates through clipped objectives and trust region constraints, effectively aligning models for dialogue systems and code generation (Wang et al., 2024b; Shojaee et al., 2023). Though joint policy-value optimization introduces computational overhead (Shao et al., 2024). GRPO (Shao et al., 2024) mitigates these limitations through trajectory-level sampling and group-wise advantage normalization. By generating multiple responses per input and standardizing rewards across trajectory groups, GRPO reduces policy update bias while maintaining efficiency. Recent implementations such as DAPO (Yu et al., 2025) and Dr. GRPO (Liu et al., 2025) further showcase GRPO’s scalability through detailed refinement of reasoning objectives, though exploration efficiency remains constrained by static and sparse reward structures (Dou et al., 2025; Gao et al., 2024). Differently, i-MENTOR enhances LLM reasoning through providing dense, stable, and efficient exploration rewards in RL optimization process. These rewards enable intrinsic motivation-guided exploration over diverse response trajectories during training, systematically improving reasoning capabilities.
@@ 5.2 Improving Reasoning Ability of LLMs
Recent advances in LLM reasoning focus on three key paradigms: Pre-training augmentation improves foundational reasoning by exposing models to curated datasets and synthetic traces. Models like Qwen (Yang et al., 2024b) and Llama (Grattafiori et al., 2024) show significant gains through domain-specific data scaling, though this requires heavy computational resources and careful data curation. Prompt engineering (Sahoo et al., 2024; Chen et al., 2023) enhances reasoning via structured inputs. Techniques like Chain-of-Thought (Yu et al., 2023; Chu et al., 2023) decompose problems into steps, while self-consistency (Wang et al., 2022) improves outputs through majority voting. Extensions such as multi-agent debates (Liang et al., 2023) and iterative refinement (Madaan et al., 2023) further extend capabilities but are limited by manual design and high computational costs. Algorithmic enhancement combines inference-time search with RL-based fine-tuning. Methods like Monte Carlo Tree Search (Browne et al., 2012) and RL approaches (e.g., PPO and GRPO) optimize reasoning policies but often overexploit fixed high-reward paths, limiting exploration of novel solutions. Unlike these methods, i-MENTOR actively incentivizes exploration of under-optimized paths through exploration rewards, enabling deeper action-space traversal.
@@ 6 Conclusion
We introduce i-MENTOR, a reinforcement learning framework that systematically addresses the challenge of sparse rewards in LLM reasoning through three methodological advances. First, our trajectory-aware exploration reward employs lightweight networks to provide dense intrinsic rewards while eliminating computational overhead and episodic length bias, enabling efficient discovery of high-quality reasoning paths. Second, we propose error-conditioned reward allocation to ensure efficient exploration on challenging samples while intrinsically stabilizing training in large action spaces, ensuring sustained exploration. Third, our advantage-preserving integration decouples exploration incentives from outcome-based policy gradients, resolving fundamental conflicts between intrinsic and extrinsic rewards. Extensive experiments demonstrate that i-MENTOR significantly enhances reasoning performance when integrated with both PPO and GRPO, establishing its versatility across distinct RL paradigms.
ETHICS STATEMENT
To the best of our knowledge, there are no ethical issues with this paper.
REPRODUCIBILITY STATEMENT
We provide our source code at https://anonymous.4open.science/r/iMENTOR-C28D to ensure reproducibility, with detailed implementation details in Appendix D.
@@ Appendix A Detailed Case Study
Through the methodological components described in Section 2.2, i-MENTOR implements a reinforcement learning framework with structured exploration guidance to enhance LLM reasoning capabilities. The complete optimization procedure, formalized in Algorithm 1, operates as follows: Given an input batch $B$ of question-response pairs $(q,o)$ for advantage estimation, we first compute the baseline outcome-based reward signals $R$ and initial advantage estimates $\hat{A}_{\text{old}}$ using conventional outcome-based reward functions and RL advantage estimation methods. Subsequently, each sample is processed through i-MENTOR’s dual networks $(f_{\theta_{P}},f_{\theta_{T}})$ to simultaneously update the policy network $f_{\theta_{P}}$ and obtain trajectory-aware exploration rewards $R^{\star}_{1}(q,o)$. These exploration rewards are then adaptively scaled through our error-conditioned reward allocation approach, yielding $R^{\star}(q,o)$ that maintains training stability across different optimization phases. Finally, the refined advantages $\hat{A}_{\text{new}}$ are computed through our advantage-preserving integration mechanism that injects $R^{\star}(q,o)$ into $\hat{A}_{\text{old}}$ without distorting the original policy gradient signals. The resulting $\hat{A}_{\text{new}}$ subsequently drives policy updates through standard RL optimization algorithms like PPO and GRPO, enabling effective LLM optimization while maintaining gradient stability.
@@
4: Obtain $R^{\star}_{1}(q,o)$ via Equation (5)
5: Obtain $R^{\star}(q,o)$ via Equation (6)
6: Add the exploration reward $R^{\star}(q,o)$ to $\hat{A}_{old}$ for a new advantage $\hat{A}_{new}$ via Equation (7)
7: return $\hat{A}_{new}$
@@ Appendix C Dataset Statistics
Appendix D Implementation Details
Appendix E Comparison with Basic Exploration Techniques
@@
Table 7: Comparison with Other Naive Exploration Methods. $\beta$” and “Temperature” indicate the KL penalty and LLM temperature coefficients.
@@
[3pt/2pt] i-MENTOR-GRPO ($\beta=0.0$, $Temp=1.0$) 0.4739
Appendix F Ablation Study
Table 8: Ablation study on Countdown-34.
Model Accuracy
GRPO 0.6711
++Trajectory-aware Exploration Rewards 0.6939
++Error-conditioned Reward Allocation 0.7065
++Advantage-Preserving Reward Integration (i-MENTOR-GRPO) 0.7132
From the table, we could conclude that: (1) Incorporating trajectory-aware exploration rewards leads to a notable improvement in reasoning performance. This demonstrates that enabling the model to explore diverse potential responses during training promotes a richer learning process. By encouraging the discovery of alternative reasoning paths rather than converging prematurely on a limited set of high-reward responses, the model acquires more nuanced reasoning capabilities, as evidenced by the marked accuracy gain. (2) Adding the error-conditioned reward allocation mechanism further enhances model performance. By applying error-conditioned reward allocation, i-MENTOR conducts efficient exploration with adaptive balance on exploration and exploitation during policy updates, ensuring stable and effective policy optimization. (3) Finally, the advantage-preserving integration ensures seamless compatibility with RL algorithms, such as PPO and GRPO. By preserving the statistical integrity of outcome-based advantage distributions while incorporating exploratory signals, i-MENTOR achieves a robust balance between exploiting known strategies and exploring new areas, thus enhancing generalization capabilities. With all three components, i-MENTOR-GRPO demonstrates the highest improvement in accuracy.
@@ Appendix G Parameter Sensitivity Analysis
In addition to Equation (5) and  (6), the detailed exploration is obtained with a more specific control on its range and attenuation rate to prioritize early-stage exploration while gradually shifting focus to exploitation, enabling stable policy convergence:
$\left\{\begin{aligned} &r(q,o)=\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2}\\ &R^{\star}_{1}(q,o)=\alpha\cdot\frac{r(q,o)-\min\limits_{(q,o)\in B}(r(q,o))}{\max\limits_{(q,o)\in B}(r(q,o))-\min\limits_{(q,o)\in B}(r(q,o))}\\ &R^{\star}_{2}(q,o)=I_{a\neq o}\cdot R^{\star}_{1}(q,o)\\ &R^{\star}(q,o)=\frac{\gamma}{\gamma+n}\cdot R^{\star}_{2}(q,o)\end{aligned}\right$. (9)
This process involves two hyperparameters: $\alpha$, which scales the maximum exploration reward intensity, and $\gamma$, which regulates its decay rate. In this section, we visualize the parameter sensitivity experiments on Countdown-4 for i-MENTOR-GRPO in Figure 6 to analyze the influence of both on the performance of i-MENTOR.